# Pool Formation in Oceanic Games: Shapley Value and Proportional Sharing

### v1

@@ 2.1. The oceanic model with atomic and non-atomic players
Remark 2.
For the remainder of the paper, we will make the assumption that $a_{i}<h$, for any atomic player $i\in N_{a}$. The main reason is that we want to focus on players who need to collaborate with other people in order to form a successful pool, as also considered in the model that initiated oceanic games (Milnor and Shapley, 1978). For players with $a_{i}>h$, given our threshold-based reward function $\rho$, intuitively it makes sense for them to break their stake into smaller players of size at most $h$ each. We essentially consider that our game begins after such a stake decomposition has taken place.
@@ 3.3.1. Games with one large player
Given Lemmas 3.13 and 3.14, we can now describe the equilibrium construction for any $n\geq l^{2}+k+1$ (for the integer values $k,l$ that were determined in Lemmas 3.13 and 3.14), and thus for any $n\geq(2h-1)^{2}+2h$. Since the values for $k$ and $l$ that we found satisfy the equilibrium conditions, all that remains to show is that we can actually partition the $n$ players into such a $(k,l)$-partition. To do this, we will have $k+1$ players, including the large player, form one Type A pool. For the remaining $n-k-1\geq(l-1)^{2}$ players, we first create as many Type C pools with exactly $l-1$ players as possible. And then, each of the remaining $(n-k-1)\mod(l-1)$ players is added to a different Type C pool, converting it to a Type B pool. This completes the proof of Theorem 3.12. ∎
@@ 3.3.2. Games with more large players
The remaining argument for completing the proof of Theorem 3.6 is in Appendix A.
@@ 5.2. Computation of Sybil strategies
The constraints above are linear but this is not always an easy problem to solve, and its complexity is dependent on the reward scheme, since this affects the form of the objective function. For simpler schemes, such as the proportional one, we can have a polynomial algorithm. In particular, under the proportional scheme, the objective function is a sum of concave functions (and hence a concave function itself), equal to $\sum_{j=1}^{m}\frac{s_{i_{j}}}{m(C_{j})+s_{i_{j}}}$. This means that the optimization problem is equivalent to minimizing a convex function subject to linear constraints. In fact, for the proportional scheme, we can have a much simpler, greedy algorithm, based on a water-filling argument (see Appendix B.1). Hence, we can conclude with the following.
@@ 7.1. Future work
Appendix A Missing proofs from Section 3
A.1. Proof of Lemma 3.9
@@ Appendix A Missing proofs from Section 3
Appendix B Missing proofs from Section 5
B.1. Proof of Theorem 5.7
@@ Appendix B Missing proofs from Section 5
Appendix C Missing proofs from Section 6
C.1. Proof of Theorem 6.3
@@ Appendix C Missing proofs from Section 6
We will show how to set values for $k$ and $l$ so that the conditions of Lemma C.1 are satisfied. From the last two conditions of Lemma C.1, we get that for the value of $k$ it must hold that $l-\sqrt{a}\leq k\leq l-\sqrt{a}+1$. Hence, there is room for $k$ to have an integer value. For $l=h+1$, which is the lowest possible value for $l$ (since it must hold that $l-1\geq h$), we get that the lowest integer value for $k$ is $\lceil h-\sqrt{a}+1\rceil$. We note that for this value of $k$ it can be verified that the first condition of Lemma C.1 is also satisfied, and also that $k+a\geq h$, which is a requirement of the lemma. Therefore, given the values of $l=h+1$ and $k=\lceil h-\sqrt{a}+1\rceil$, and if the number of small players is sufficiently large, we can construct a $(k,l)$-partition in the same way as in the last part in the proof of Theorem 3.12. Recalling that $n_{b}$ is the number of the big players, the total number of pools created for the values of $k$ and $l$ we use is at least: $n_{b}+\lfloor\frac{n-n_{b}-n_{b}\cdot(h-\sqrt{a}+2)}{h+1}\rfloor=\lfloor\frac{n_{s}+n_{b}\cdot(\sqrt{a}-1)}{h+1}\rfloor$.
@@
From the previous theorem, we can actually extract a constant upper bound of 2 on PoS. Under the conditions of Theorem C.2, we show that there always exists an equilibrium with $l-2=h$ and $k=\lceil h-\sqrt{a}+1\rceil\leq h-\sqrt{a}+2\leq h+2$. To establish that we have an equilibrium for these choices of $k$ and $l$, it suffices to show that for $l=h+2$ and $k=\lceil h-\sqrt{a}+1\rceil$, the conditions of Lemma C.1 are met. Firstly, for $l=h+2$ the first condition is met, since $h+2>k+1$. Moreover, for the value of $k$ we have that $\lceil h-\sqrt{a}+1\rceil\geq h-\sqrt{a}+1=l-\sqrt{a}-$ and also $\lceil h-\sqrt{a}+1\rceil\leq h-\sqrt{a}+2=l-\sqrt{a}$. Hence, the last two conditions are also satisfied.
@@
At the same time, we can also have a matching lower bound on PoS, making the previous upper bound tight. A simple instance in which the scheme achieves Price of Stability equal to 2 is the following: there is one player with stake $a$ and $2h-a$ players with stake 1. The optimal formation consists of 2 pools with total stake exactly $h$. From Lemma C.1, in order for a formation with $l=h$ to be an equilibrium, it must hold that $k\geq l-\sqrt{a}-1=h-\sqrt{a}-1\geq h-a$, for any $a\geq 3$ and hence there are not enough small players for a Nash equilibrium with 2 pools to exist. Thus only the grand coalition is an equilibrium, and hence PoS = 2.
@@
C.3. Proof of Theorem 6.5
In order to show this, we use the $(k,l)$-equilibrium identified in Appendix C.2 with $k=\lceil h-\sqrt{a}-1\rceil$. The reward of each large agent then is $\frac{\sqrt{a}}{\sqrt{a}+h-\lceil\sqrt{a}\rceil-1}$. Consider the scenario where a large agent $i$ splits her stake into single unit amounts and commits them to other Type A pools (we consider an instance where there are at least $a+1$ large players). The total reward she receives in that case is: $\frac{a}{\sqrt{a}+h-\lceil\sqrt{a}\rceil}$. We can now check that $\frac{a}{\sqrt{a}+h-\lceil\sqrt{a}\rceil}>\frac{\sqrt{a}}{\sqrt{a}+h-\lceil\sqrt{a}\rceil-1}$, for any $2\leq a<h$, thus the player receives more than committing to a single pool.

### v2

@@ 2.1. The oceanic model with atomic and non-atomic players
Remark 1.
For the remainder of the paper, we make the assumption that $a_{i}<h$, for any atomic player $i\in N_{a}$. The main reason is that, as also considered in the initial model of oceanic games by (Milnor and Shapley, 1978), we want to focus on players who need to collaborate with other people in order to form a successful pool. If there exist players with $a_{i}>h$, given our threshold-based reward function $\rho$, intuitively it makes sense for them to break their stake and run their own pools with avatars of size $h$ each, and participate in our game with the remainder $a_{i}~{}mod~{}h$. For a further discussion on this, see Section A in the Appendix.
@@ 3.3.1. Games with one large player
Given Lemmas 3.12 and 3.13, we can now describe the equilibrium construction for any $n\geq l^{2}+k+1$ (for the integer values $k,l$ that were determined in Lemmas 3.12 and 3.13), and thus for any $n\geq(2h-1)^{2}+2h$. Since the values for $k$ and $l$ that we found satisfy the equilibrium conditions, all that remains to show is that we can actually partition the $n$ players into such a $(k,l)$-partition. To do this, we will have $k+1$ players, including the large player, form one Type A pool. For the remaining $n-k-1\geq(l-1)^{2}$ players, we first create as many Type C pools with exactly $l-1$ players as possible. And then, each of the remaining $(n-k-1)\mod(l-1)$ players is added to a different Type C pool, converting it to a Type B pool. This completes the proof of Theorem 3.11. ∎
@@ 3.3.2. Games with more large players
The remaining argument for completing the proof of Theorem 3.6 is in Appendix B.
@@ 5.2. Computation of Sybil strategies
The constraints above are linear but this is not always an easy problem to solve, and its complexity is dependent on the reward scheme, since this affects the form of the objective function. For simpler schemes, such as the proportional one, we can have a polynomial algorithm. In particular, under the proportional scheme, the objective function is a sum of concave functions (and hence a concave function itself), equal to $\sum_{j=1}^{m}\frac{s_{i_{j}}}{m(C_{j})+s_{i_{j}}}$. This means that the optimization problem is equivalent to minimizing a convex function subject to linear constraints. In fact, for the proportional scheme, we can have a much simpler, greedy algorithm, based on a water-filling argument (see Appendix C.1). Hence, we can conclude with the following.
@@ 7.1. Future work
Appendix A Discussion related to Remark 1
In this section, we revisit the assumption we have made that $a_{i}<h$. If we have very large players with $a_{i}>h$, obviously one choice for them is that they could split their stake into portions of size $h$ and run their own individual pools, which will not get involved in the pool formation game. Hence they would participate in pools with other players only with their remainder, which is $a_{i}~{}mod~{}h$. Given our equilibrium analysis, we provide evidence below that this is a meaningful action for them, in the sense that splitting their stake in portions smaller than $h$ (say $h/m$ for some $m>1$) and participate with multiple avatars into our pool formation game does not yield better outcomes for them.
Theorem A.1.
Consider a stake distribution where there exists a single player $i\in N_{a}$ with $a_{i}=\lambda h$ and for every $j\neq i$, $a_{i}<h$. Suppose that player $i$ is considering splitting her stake into portions of size $h/m$ and forming pools with other players, where $m>1$. Then, for the Shapley scheme, and the equilibria that arise under this splitting, as identified in Section 3, player $i$ is not better off, compared to the scenario where she splits her stake into $\lambda$ portions of size exactly $h$ each and runs her own individual pools. The same also holds for any equilibrium under the proportional scheme.
Proof.
If $i$ splits her stake into portions of size $h$, and runs her own pools, her payoff will be exactly equal to $\lambda$. We compare this against a uniform split into $\lambda m$ portions of size $h/m$ each, where $m>1$. Hence player $i$ participates with $\lambda m$ identities in the game.
Consider the Shapley scheme. For any equilibrium in the form described by the conditions of Lemma 3.5, each such avatar will form a pool with small players. By Lemma 3.2, the reward of each avatar will be equal to its Shapley value in a pool consisting of herself and a volume of mass $k_{i}$ for some $k_{i}\leq h$, which is equal to $\frac{k_{i}-h+h/m}{k_{i}}$. Hence, the total reward of player $i$, after summing up the reward of all her avatars will be equal to:
$\lambda m\cdot\frac{k_{i}-h+h/m}{k_{i}}$
If we now check for what values of $m$ this total reward is no more than $\lambda$, this is equivalent to:
$m\cdot\frac{k_{i}-h+h/m}{k_{i}}\leq 1$
By simplifying the above expression, we get that this is equivalent to $k_{i}\leq h$, which is always true for these equilibria by Lemma 3.5. Hence, $i$ cannot receive a higher reward by splitting her stake uniformly into portions of size strictly less than $h$.
Finally, regarding the proportional scheme, it is very easy to verify that a player cannot become better off by splitting her stake into portions of size $h/m$ each, for $m>1$. ∎
Appendix B Missing proofs from Section 3
B.1. Proof of Lemma 3.8
@@ Appendix B Missing proofs from Section 3
B.6. Proof of Theorem 3.6
Lemma B.2.
Proof.
and the second one is
$(k+2)(k-h+a+1)\geq(h-a)(h-a+1)$
Appendix C Missing proofs from Section 5
C.1. Proof of Theorem 5.7
@@ Appendix C Missing proofs from Section 5
Appendix D Missing proofs from Section 6
D.1. Proof of Theorem 6.3
@@ Appendix D Missing proofs from Section 6
We will show how to set values for $k$ and $l$ so that the conditions of Lemma D.1 are satisfied. From the last two conditions of Lemma D.1, we get that for the value of $k$ it must hold that $l-\sqrt{a}\leq k\leq l-\sqrt{a}+1$. Hence, there is room for $k$ to have an integer value. For $l=h+1$, which is the lowest possible value for $l$ (since it must hold that $l-1\geq h$), we get that the lowest integer value for $k$ is $\lceil h-\sqrt{a}+1\rceil$. We note that for this value of $k$ it can be verified that the first condition of Lemma D.1 is also satisfied, and also that $k+a\geq h$, which is a requirement of the lemma. Therefore, given the values of $l=h+1$ and $k=\lceil h-\sqrt{a}+1\rceil$, and if the number of small players is sufficiently large, we can construct a $(k,l)$-partition in the same way as in the last part in the proof of Theorem 3.11. Recalling that $n_{b}$ is the number of the big players, the total number of pools created for the values of $k$ and $l$ we use is at least: $n_{b}+\lfloor\frac{n-n_{b}-n_{b}\cdot(h-\sqrt{a}+2)}{h+1}\rfloor=\lfloor\frac{n_{s}+n_{b}\cdot(\sqrt{a}-1)}{h+1}\rfloor$.
@@
From the previous theorem, we can actually extract a constant upper bound of 2 on PoS. Under the conditions of Theorem D.2, we show that there always exists an equilibrium with $l-2=h$ and $k=\lceil h-\sqrt{a}+1\rceil\leq h-\sqrt{a}+2\leq h+2$. To establish that we have an equilibrium for these choices of $k$ and $l$, it suffices to show that for $l=h+2$ and $k=\lceil h-\sqrt{a}+1\rceil$, the conditions of Lemma D.1 are met. Firstly, for $l=h+2$ the first condition is met, since $h+2>k+1$. Moreover, for the value of $k$ we have that $\lceil h-\sqrt{a}+1\rceil\geq h-\sqrt{a}+1=l-\sqrt{a}-$ and also $\lceil h-\sqrt{a}+1\rceil\leq h-\sqrt{a}+2=l-\sqrt{a}$. Hence, the last two conditions are also satisfied.
@@
At the same time, we can also have a matching lower bound on PoS, making the previous upper bound tight. A simple instance in which the scheme achieves Price of Stability equal to 2 is the following: there is one player with stake $a$ and $2h-a$ players with stake 1. The optimal formation consists of 2 pools with total stake exactly $h$. From Lemma D.1, in order for a formation with $l=h$ to be an equilibrium, it must hold that $k\geq l-\sqrt{a}-1=h-\sqrt{a}-1\geq h-a$, for any $a\geq 3$ and hence there are not enough small players for a Nash equilibrium with 2 pools to exist. Thus only the grand coalition is an equilibrium, and hence PoS = 2.
@@
D.3. Proof of Theorem 6.5
In order to show this, we use the $(k,l)$-equilibrium identified in Appendix D.2 with $k=\lceil h-\sqrt{a}-1\rceil$. The reward of each large agent then is $\frac{\sqrt{a}}{\sqrt{a}+h-\lceil\sqrt{a}\rceil-1}$. Consider the scenario where a large agent $i$ splits her stake into single unit amounts and commits them to other Type A pools (we consider an instance where there are at least $a+1$ large players). The total reward she receives in that case is: $\frac{a}{\sqrt{a}+h-\lceil\sqrt{a}\rceil}$. We can now check that $\frac{a}{\sqrt{a}+h-\lceil\sqrt{a}\rceil}>\frac{\sqrt{a}}{\sqrt{a}+h-\lceil\sqrt{a}\rceil-1}$, for any $2\leq a<h$, thus the player receives more than committing to a single pool.